The following financial statement data for years ending December 31 for Foodworks Company are shown below. 2016 2015 Sales $838,500 $544,500 Total assets: Beginning of year 560,000 430,000 End of year 730,000 560,000 a. Determine the ratio of sales to assets for 2016 and 2015. Round your answers to one decimal place. 2016 2015 Ratio of sales to assets b. Does the change in the ratio of sales to assets from 2015 to 2016 indicate a favorable or an unfavorable trend?
Answer #1

Calculate ratio of sales of assets

2016 2015
Sales 838500 544500
Average assets (560000+730000/2) = 645000 (430000+560000/2) = 495000
Ratio of sales to assets 838500/645000 = 1.3 544500/495000 = 1.10

Favorable trend